Deposits are made at the end of years 1 to 7 into an account paying 5.8%/year compounded annually. The deposits start at 5000 and increase by $1000 each year. How much will be in the account after the last deposit?
Answer #1
Year Deposit
1 5000
2 (5000+1000)=6000
3 (6000+1000)=7000
4 (7000+1000)=8000
5 (8000+1000)=9000
6 (9000+1000)=10000
7 (10000+1000)=11000

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=5000*(1.058)^6+6000*(1.058)^5+7000*(1.058)^4+8000*(1.058)^3+9000*(1.058)^2+10,000*(1.058)^1+11000

which is equal to

=$64865.97(Approx).